I can't figure out the formula for each one?Home Insert Draw Page Layout Formulas Data Review View 2 E E N DO E Get Add-ins G?- c rv - the slicer e Timeline Link New ComB G H Problem 2: The ideal gas law equation relates the pressure, temperature and volume of a gas. It states: PV = nRT whereHome Insert DawPage Layout Formulas Data Review View Get Add-ins L Slicer Timeline Pivot Table Recommended Table Pivot TablesE23 fix ABCDEFGHIJKLMNOPI Problem 5: The time dilation effect of general relativity can be described by: The solution to thisProblem 6: In trigonometry, the sine of the sum of angles follow the identity: sin( + ) = sin cos + cos sin Using the built-i
Answer #1

Since the actual equations are already given with the question, I assume that what you need is the code that has to be typed in excel to get the answer.

Any cell in excel is identified by the letter on the top of the row and the number on the left of the column.

1) In the cell D15, type the following.

=2*(D13-D9)/(D10*D10+2*D11*D12)

and press enter. You will get the answer.(dont miss the = sign or any character.)

2) In the cell D12, type the following.

=(D13*D14)/(D15*D16)

and press enter. You will get the answer.

3) In the cell D17, type the following.

=SQRT(2*(D15-D11)/D12-D13*D14)

and press enter. You will get the answer.

4) In the cell D16, type the following.

=D14*SQRT(1-(D12/D13)*(D12/D13))

and press enter. You will get the answer.

5) In the cell D10, type the following.

=SIN(D8+D9)

and press enter.

In the cell D11, type the following.

=SIN(D8)*COS(D9)

and press enter.

In the cell D12, type the following.

=COS(D8)*SIN(D9)

and press enter.

In the cell D13, type the following.

=D11+D12

and press enter.

If the values in D10 and D13 are equal, type 'yes' in D15